A particle of mass 6.5 times 10^-8 kg and charge +
Answer #1

Mass of particle, m = 6.5 \times 10^{-8} kg

Charge of particle, q =9.7 \times 10^{-6} C

Existing Magnetic Field, B= 4.4 T

If the particle travels a semi-circular path, then distance travelled by it is

  s= \pi r

where radius of the path is given by

r= \frac{mv}{qB}

where v= speed of particle

Then time spent will be t= \frac{s}{v}

  t= \frac{\pi r}{v}= \frac{\pi mv}{vqB}= \frac{\pi m}{qB}

using given values in above,

  t= \frac{3.14 \times 6.5 \times 10^{-8}}{9.7 \times 10^{-6}\times 4.4}= 4.78 \times 10^{-3}s(ANS)